# Maximizing Relation Extraction Potential: A Data-Centric Study to Unveil Challenges and Opportunities

### v1

@@ V-C Long-tail Data Distribution
The problem of long-tail relations that arise from datasets with imbalanced classes proved to be a significant disruptor for supervised relation extractors. This problem is a critical bottleneck in the relation extraction pipeline, as supervised and joint extractors cannot handle relation classes with negligible data samples. However, adding more data to balance out the long-tail problem is not always possible. With the recent dominance of LLMs in the NLP literature in low-resource scenarios, it is imperative to develop better LLM-based solutions for relation extractions that are robust to some of the issues mentioned in this study. The alleviation of this issue can prove to be critical in domains where there is not enough data, such as biomedical and personality prediction.
@@ V-D Fine-grained Relation Distribution
The findings in this study highlight the inefficiency of the LLMs in extracting relations from fine-grained datasets with few-shot learning. Relation extraction on fine-grained datasets that involve numerous labels is a challenging task as the possibility of similar relation types increases as more relations are introduced. However, these datasets pose an even more challenging use case for LLMs. It is impractical to incorporate the verbalization of the full label space in the prompt due to the restricted token limit. Furthermore, it is difficult to add demonstrations of each relation type in the prompt as the number of relations increases. Thus, future work is needed to improve LLMs at tackling fine-grained relation distributions. Possible directions for future work could be the development of better-prompting strategies that can provide concise knowledge of the label space. Research in this direction would significantly impact relationship extraction from complex domains such as finance and business.
@@ VI Conclusion
Appendix A Datasets
@@ Appendix C Results
[t] Datasets Algorithms K-shots 1 5 10 20 30 FewRel RBERT 5.43$\pm$0.24 30.03$\pm$0.12 56.02$\pm$0.9 69.45$\pm$0.98 77.83$\pm$1.07 Roberta_base 1.7$\pm$0.12 2.63$\pm$0.87 11.84$\pm$1.07 36.79$\pm$2.27 50.88$\pm$3.37 DeepKE1 35.67$\pm$1.17 - - - - GPT-RE1 46.9$\pm$0.52 54.0$\pm$0.29 56.84$\pm$0.21 57.47$\pm$0.27 58.36$\pm$0.19 GenPT/Bart 43.47$\pm$1.03 69.56$\pm$0.57 77.29$\pm$0.57 82.61$\pm$0.36 84.56$\pm$0.86 GenPT/T5 10.15$\pm$1.23 55.34$\pm$0.81 73.36$\pm$0.42 81.34$\pm$0.77 84.08$\pm$0.56 GenPT/roberta 31.27$\pm$1.09 72.52$\pm$0.9 79.84$\pm$0.65 82.19$\pm$1.0 84.63$\pm$0.42 KnowPrompt 7.45$\pm$0.19 37.57$\pm$0.68 55.93$\pm$1.36 73.39$\pm$2.0 80.22$\pm$0.61 NYT10 RBERT 1.77$\pm$0.17 12.58$\pm$0.66 21.29$\pm$2.23 40.07$\pm$1.47 48.15$\pm$2.85 Roberta_base 7.82$\pm$1.5 1.64$\pm$2.1 7.77$\pm$2.54 5.69$\pm$2.07 10.51$\pm$2.94 DeepKE1 50.66$\pm$0.97 - - - - GPT-RE1 72.48$\pm$1.17 79.44$\pm$0.15 77.16$\pm$0.2 75.66$\pm$0.58 74.73$\pm$0.7 GenPT/Bart 3.04$\pm$5.08 21.26$\pm$2.34 31.34$\pm$2.18 41.39$\pm$1.08 44.4$\pm$4.55 GenPT/T5 6.61$\pm$0.46 19.99$\pm$0.3 35.28$\pm$0.24 43.84$\pm$1.57 47.81$\pm$1.72 GenPT/roberta 3.23$\pm$0.48 26.69$\pm$1.98 37.18$\pm$1.84 47.78$\pm$2.41 50.16$\pm$2.19 KnowPrompt 0.31$\pm$0.01 2.14$\pm$0.48 11.61$\pm$1.1 23.39$\pm$2.23 32.37$\pm$2.93 WebNLG RBERT 9.98$\pm$0.6 40.47$\pm$0.32 56.59$\pm$1.22 63.65$\pm$0.9 69.97$\pm$1.03 Roberta_base 0.83$\pm$0.26 5.23$\pm$0.5 18.98$\pm$2.16 38.94$\pm$6.15 50.47$\pm$0.98 DeepKE1 - - - - - GPT-RE1 28.57$\pm$0.58 27.99$\pm$0.46 29.23$\pm$0.36 28.31$\pm$1.33 27.89$\pm$2.21 GenPT/Bart 45.56$\pm$2.36 74.19$\pm$0.44 79.13$\pm$0.45 85.26$\pm$0.38 87.42$\pm$1.74 GenPT/T5 41.49$\pm$1.55 66.98$\pm$0.77 77.82$\pm$0.62 84.83$\pm$0.92 87.6$\pm$0.92 GenPT/roberta 51.27$\pm$1.74 72.48$\pm$1.81 80.6$\pm$1.19 86.11$\pm$1.42 87.85$\pm$0.49 KnowPrompt 3.05$\pm$0.72 48.63$\pm$1.25 60.58$\pm$1.44 72.06$\pm$1.84 81.62$\pm$0.61 crossRE RBERT 7.85$\pm$0.4 17.96$\pm$1.57 21.47$\pm$1.23 28.46$\pm$1.11 38.28$\pm$2.95 Roberta_base 4.36$\pm$0.9 6.97$\pm$1.25 9.33$\pm$3.21 6.5$\pm$1.74 20.9$\pm$4.32 DeepKE1 20.04$\pm$0.93 - - - - GPT-RE1 30.99$\pm$1.16 41.1$\pm$0.69 45.16$\pm$0.16 46.69$\pm$0.67 48.75$\pm$0.74 GenPT/Bart 5.7$\pm$0.75 32.39$\pm$3.99 37.68$\pm$1.28 45.12$\pm$2.51 53.18$\pm$2.09 GenPT/T5 5.75$\pm$0.12 10.51$\pm$0.86 23.31$\pm$5.13 44.22$\pm$1.7 50.16$\pm$0.65 GenPT/roberta 4.89$\pm$0.49 34.14$\pm$2.75 39.12$\pm$3.49 50.48$\pm$2.22 58.8$\pm$1.56 KnowPrompt 6.04$\pm$0.04 7.9$\pm$2.1 4.5$\pm$0.65 22.96$\pm$3.19 28.64$\pm$4.46 SemEval2 RBERT 9.67$\pm$0.52 16.42$\pm$0.38 25.75$\pm$0.89 34.63$\pm$2.36 44.04$\pm$2.34 Roberta_base 7.21$\pm$0.81 10.41$\pm$1.26 7.23$\pm$3.58 14.7$\pm$1.91 13.45$\pm$2.01 DeepKE1 45.74$\pm$1.98 - - - - GPT-RE1 51.03$\pm$0.86 62.78$\pm$0.2 61.33$\pm$0.55 60.54$\pm$0.67 60.67$\pm$0.35 GenPT/Bart 8.04$\pm$0.99 33.82$\pm$2.15 46.51$\pm$1.84 60.66$\pm$2.37 47.12$\pm$32.94 GenPT/T5 16.58$\pm$1.0 13.67$\pm$1.99 33.96$\pm$1.94 48.35$\pm$3.4 58.68$\pm$1.33 GenPT/roberta 14.94$\pm$1.82 29.75$\pm$1.7 47.93$\pm$2.32 63.4$\pm$1.29 70.53$\pm$1.29 KnowPrompt 10.72$\pm$0.09 10.73$\pm$0.1 9.96$\pm$1.18 24.33$\pm$5.64 37.76$\pm$8.43 TACRED RBERT 14.19$\pm$3.37 4.84$\pm$1.46 9.19$\pm$0.98 14.56$\pm$1.56 25.58$\pm$2.83 Roberta_base 1.92$\pm$0.76 1.34$\pm$3.13 1.33$\pm$0.39 6.58$\pm$0.63 11.91$\pm$1.45 DeepKE1 68.8$\pm$4.86 - - - - GPT-RE1 3.47$\pm$0.2 6.96$\pm$0.23 8.84$\pm$0.04 10.26$\pm$0.15 10.79$\pm$0.07 GenPT/Bart 22.72$\pm$0.26 26.69$\pm$1.65 27.12$\pm$4.27 28.97$\pm$5.1 31.92$\pm$1.1 GenPT/T5 24.72$\pm$1.25 21.65$\pm$1.04 19.8$\pm$1.96 25.96$\pm$5.66 28.0$\pm$3.52 GenPT/roberta 20.17$\pm$1.24 20.58$\pm$2.78 19.95$\pm$0.43 18.62$\pm$1.44 18.36$\pm$1.94 KnowPrompt 0.52$\pm$0.03 3.89$\pm$0.33 11.73$\pm$0.19 20.32$\pm$1.27 28.02$\pm$2.02 RETACRED RBERT 5.7$\pm$1.32 9.25$\pm$0.44 16.58$\pm$0.37 29.71$\pm$2.02 36.04$\pm$2.01 Roberta_base 0.17$\pm$0.21 1.67$\pm$0.34 12.41$\pm$3.19 14.71$\pm$1.61 19.65$\pm$4.62 DeepKE1 57.99$\pm$1.49 - - - - GPT-RE1 13.83$\pm$0.25 27.66$\pm$0.5 30.63$\pm$0.47 32.67$\pm$0.12 32.8$\pm$0.36 GenPT/Bart 29.34$\pm$0.36 41.54$\pm$1.39 47.93$\pm$2.43 53.56$\pm$1.02 54.44$\pm$5.14 GenPT/T5 27.75$\pm$0.96 32.06$\pm$1.24 41.95$\pm$1.28 48.8$\pm$3.3 53.09$\pm$5.82 GenPT/roberta 17.52$\pm$1.02 39.33$\pm$1.23 49.33$\pm$1.9 56.22$\pm$2.15 59.07$\pm$0.95 KnowPrompt 0.61$\pm$0.02 8.74$\pm$0.16 18.77$\pm$0.54 29.03$\pm$0.18 39.94$\pm$3.13
@@
Method FewRel NYT10 WebNLG crossRE RETACRED SemEval TACRED
RIFRE 35.93$\pm$0.53 72.88$\pm$0.5 90.16$\pm$0.41 0.72$\pm$0.28 - 64.27$\pm$0.39 -
SPN4RE 36.67$\pm$0.12 72.66$\pm$0.19 88.38$\pm$0.21 31.0$\pm$0.79 20.66$\pm$1.39 60.56$\pm$0.68 19.81$\pm$0.83
TDEER 33.98$\pm$0.62 71.86$\pm$0.25 70.81$\pm$16.41 31.88$\pm$4.04 16.46$\pm$0.84 58.79$\pm$1.04 15.1$\pm$0.8
UniRel 39.53$\pm$0.76 67.72$\pm$1.61 89.75$\pm$0.27 43.5$\pm$0.3 - 61.56$\pm$0.68 -

### v2

@@ V-C Long-tail Data Distribution
The problem of long-tail relations that arise from datasets with imbalanced classes proved to be a significant disruptor for supervised relation extractors. This problem is a critical bottleneck in the relation extraction pipeline, as supervised and joint extractors cannot handle relation classes with negligible data samples. A probable solution to this problem could be to harness the generative nature of LLMs and use them as synthetic data generators to balance out the long tail of the distribution [84]. If the creation of novel data is not feasible, then the potential of LLM-based relation extractors should be explored to tackle this complex use case. The low resource inference capabilities of such models afforded by pre-training on internet-scale datasets make them a natural solution to this problem. The alleviation of issues arising from the long-tail distribution can prove to be critical in domains where the availability of labeled data is scarce, such as biomedical and personality prediction.
@@ V-D Fine-grained Relation Distribution
The findings in this study highlight the inefficiency of the LLMs in extracting relations from fine-grained datasets with few-shot learning. Relation extraction on fine-grained datasets that involve numerous labels is a challenging task as the possibility of similar relation types increases as more relations are introduced. However, these datasets pose an even more challenging use case for LLMs. It is impractical to incorporate the verbalization of the full label space in the prompt due to the restricted token limit. Furthermore, it is difficult to add demonstrations of each relation type in the prompt as the number of relations increases. Thus, future work is needed to improve LLMs at tackling fine-grained relation distributions. A possible direction could be the development of better-prompting strategies that can provide concise knowledge of the label space. For example, KBs contain essential knowledge regarding numerous entity and relation types; incorporating such knowledge sources in place of label verbalizations and demonstrations might help provide a comprehensive image of the label space and aid the extraction process [85]. Research in developing robust fine-grained relation extractors would significantly impact relationship extraction from complex domains such as finance and business.
@@ VI Conclusion
Appendix A Additional Details
A-A Datasets
@@ Appendix B Results
[t] Method Dataset k 1 5 10 20 30 RBERT FewRel 5.11 30.18 55.2 69.88 77.83 NYT10 2.02 12.93 23.53 40.52 48.16 WebNLG 10.28 40.19 57.26 64.11 69.97 CrossRE 7.25 16.33 20.86 27.75 38.28 RETACRED 2.05 8.86 18.67 34.65 42.63 SemEval2 10.55 16.4 26.74 39.5 48.19 TACRED 1.97 4.7 9.84 17.78 25.01 Roberta_base FewRel 1.52 3.68 11.14 36.46 50.88 NYT10 6 3.27 5.31 6.93 10.51 WebNLG 0.44 4.67 18.07 42.18 50.47 CrossRE 3.26 7.82 6.72 5.42 20.9 RETACRED 0.61 1.74 12.42 20.39 27.65 SemEval2 9.13 9.59 4.7 14.58 12.61 TACRED 1.26 1.79 1.95 10.56 19.61 UnleashLLM1 FewRel 35.67 - - - - NYT10 50.66 - - - - WebNLG - - - - - CrossRE 20.04 - - - - RETACRED 40.4 - - - - SemEval2 50.51 - - - - TACRED 44.69 - - - - GPT-RE1 FewRel 47.52 54.1 57.02 57.64 58.36 NYT10 74.26 79.49 77.29 76.03 74.73 WebNLG 27.79 27.78 28.9 27.47 27.89 CrossRE 32.71 41.99 45.07 47.07 48.75 RETACRED 51.91 69.5 71.71 74.13 74.96 SemEval2 68.35 75.04 73.2 72.27 71.8 TACRED 36.52 45.56 47.91 50.71 53.4 GenPT/Bart FewRel 44.87 70.29 76.73 82.44 84.56 NYT10 10.65 24.23 33.42 40.69 44.4 WebNLG 49.1 74.14 78.88 85.11 87.41 CrossRE 4.58 27.34 38.81 45.92 53.18 RETACRED 33.7 51.91 54.53 60.84 60.37 SemEval2 10.31 36.79 52.41 65.04 51.01 TACRED 20.54 30.3 32.73 34.79 35.03 GenPT/T5 FewRel 11.25 54.55 73.49 81.8 84.08 NYT10 6.15 20.1 35.33 44.83 47.81 WebNLG 43.74 66.6 78.32 85.36 87.6 CrossRE 5.55 10.11 21.41 43.13 50.15 RETACRED 35.06 45.09 53.56 57.91 60.93 SemEval2 6.87 16.48 38.4 53.75 63.59 TACRED 18.35 25.7 29.64 33.49 33.86 GenPT/roberta FewRel 32.17 73.66 79.41 82.82 84.63 NYT10 3.77 29.06 35.33 48.71 50.16 WebNLG 53.77 71.02 79.57 86.04 87.85 CrossRE 5.31 31.39 42.59 51.75 58.8 RETACRED 16.73 52.16 56.9 61.5 63.89 SemEval2 3.88 34.33 53.57 68.91 75.6 TACRED 18.86 25.34 21.41 22.18 20.01 KnowPrompt FewRel 7.41 37.11 55.76 73.3 80.22 NYT10 0.29 1.55 10.94 24.75 32.37 WebNLG 4.05 47.6 60.25 71.96 81.62 CrossRE 5.98 5.23 3.91 21.33 28.64 RETACRED 0.83 9.07 23.91 37.62 47.76 SemEval2 11.8 11.8 12 23.89 41.1 TACRED 0.88 5.38 17.08 27.58 30.87
@@
Method FewRel NYT10 WebNLG crossRE SemEval TACRED RETACRED
RIFRE 35.93$\pm$0.53 72.88$\pm$0.5 90.16$\pm$0.41 0.72$\pm$0.28 64.27$\pm$0.39 - -
SPN4RE 36.67$\pm$0.12 72.66$\pm$0.19 88.38$\pm$0.21 31.0$\pm$0.79 60.56$\pm$0.68 19.81$\pm$0.83 20.66$\pm$1.39
TDEER 33.98$\pm$0.62 71.86$\pm$0.25 70.81$\pm$16.41 31.88$\pm$4.04 58.79$\pm$1.04 15.1$\pm$0.8 16.46$\pm$0.84
UniRel 39.53$\pm$0.76 67.72$\pm$1.61 89.75$\pm$0.27 43.5$\pm$0.3 61.56$\pm$0.68 - -